# Word Order in English-Japanese Simultaneous Interpretation: Analyses and Evaluation using Chunk-wise Monotonic Translation

### v1

@@ Abstract
This paper analyzes the features of monotonic translations, which follow the word order of the source language, in simultaneous interpreting (SI). The word order differences are one of the biggest challenges in SI, especially for language pairs with significant structural differences like English and Japanese. We analyzed the characteristics of monotonic translations using the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset and found some grammatical structures that make monotonic translation difficult in English-Japanese SI. We further investigated the features of monotonic translations through evaluating the output from the existing speech translation (ST) and simultaneous speech translation (simulST) models on NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as on existing test sets. The results suggest that the existing SI-based test set underestimates the model performance. We also found that the monotonic-translation-based dataset would better evaluate simulST models, while using an offline-based test set for evaluating simulST models underestimates the model performance.
@@ 1 Introduction
Simultaneous Interpreting (SI) is the task of translating speech from a source language into a target language in real time. SI is a cognitively demanding task, and human simultaneous interpreters are known to employ strategies such as segmentation, summarization, and generalization He et al. (2016). Maintaining the word order in a source language is also an important strategy, especially for language pairs whose word order is different (e.g., English and Japanese), to shorten delay and reduce cognitive load. Because of these features, SI sentences are different from offline translation sentences, but most of the studies in automatic SI Oda et al. (2014); Ma et al. (2019); Liu et al. (2020); Papi et al. (2023) have been used offline translation corpora (e.g., MuST-C; Di Gangi et al., 2019) for both training and evaluation of the models due to the limited amount of simultaneous interpretation corpora (SICs).
For the English-Japanese language pair, several SICs have been constructed Toyama et al. (2004); Shimizu et al. (2014); Matsushita et al. (2020); Doi et al. (2021). Based on NAIST Simultaneous Interpretation Corpus (NAIST-SIC; Doi et al., 2021), Zhao et al. (2024)111The dataset was released in 2023 (see the version 3 of the paper). created an automatically-aligned parallel SI dataset, named NAIST-SIC-Aligned. Since the sentences in the dataset are aligned at the sentence level, they can be used for model training. Actually, Ko et al. (2023) and Zhao et al. (2024) trained SI models using SI data from NAIST-SIC-Aligned. The model performance was evaluated through automatic evaluation metrics such as BLEU Papineni et al. (2002) using a small test set curated based on SI sentences generated by professional human simultaneous interpreters.
The scores reported in Ko et al. (2023) and Zhao et al. (2024) were relatively low but the test set used in the two studies might have underestimated the model performance. Since human simultaneous interpreters use strategies such as summarization and generalization, phrases that do not affect the main idea are not necessarily translated into the target language. If an SI model generates translation for the phrases that a human interpreter did not, the output sentence might not be evaluated properly, even when it is correct translation.
@@
Fukuda et al. (2024) pointed out the difficulty for SI models to learn which phrases in source speech are less important and called for constructing SI models that only employ the strategy of maintaining the word order in a source language. As the first step, they created NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset222https://dsc-nlp.naist.jp/data/NAIST-SIC/Aligned-Chunk_Mono-EJ/. The source sentences in the test set used in Ko et al. (2023) were automatically segmented into chunks, and each chunk was translated in a way that did not include the content of the subsequent chunks. Unlike in SI sentences by human interpreters, where not all information in the source sentences is translated, chunk-wise monotonic translation (CMT) was developed so that all information is translated (Table 1)333Precisely, omission to maintain the fluency of the sentence was allowed. See Section 3.1 for the details about the dataset.. Fukuda et al. (2024) have investigated the quality of CMT through human evaluation, but have not analyzed the characteristics of it. In addition, they have not conducted any evaluation experiments in which model outputs are evaluated on the CMT Evaluation dataset.
@@
In this paper, we analyze CMT sentences qualitatively and quantitatively. In the process of developing CMT, it was allowed to repeat, defer, and omit phrases in the source sentences to keep the translation fluent. We suppose that there are factors (e.g., syntactic structures) that prevent monotonic translation if phrases were repeated, deferred, or omitted in CMT sentences since the sentences were translated without time constraints. In addition, we evaluate the output from (1) a speech translation (ST) model trained on offline data, (2) a simultaneous speech translation (simulST) model trained on offline data, and (3) a simulST model trained on both offline and SI data. The contributions of this paper are as follows:
@@
We analyze CMT sentences and show that they tend to be longer than offline translation primarily because of repetition.
@@
We investigate what makes the phrases in source sentences to be repeated, deferred, and omitted, and show that most cases occur because of particular grammatical structures. When a phrase in a chunk is a dependent of a phrase in the preceding chunk, the head phrase is typically repeated or deferred.
@@
We evaluate the output from three different models on NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset: (1) an ST model trained on offline data, (2) a simulST model trained on offline data, and (3) a simulST model trained on both offline and SI data. We show that the existing SI-based test set Ko et al. (2023); Zhao et al. (2024) underestimates the model performance. The results also show that the CMT evaluation dataset would better evaluate simulST models while using an offline-based test set for evaluating simulST models underestimates the model performance.
@@ 2.1 Simultaneous Interpretation Corpora
SICs are valuable resources both for developing automatic SI models and for analyzing the characteristics of SI. For the English-Japanese language pair, several SICs are publicly available Toyama et al. (2004); Shimizu et al. (2014); Matsushita et al. (2020); Doi et al. (2021) although the amount of such corpora is very limited compared to offline translation corpora.
Using these corpora, SI sentences have been analyzed from various perspectives, such as strategies and interpreting patterns used by interpreters, latency, translation quality, and word order Tohyama and Matsubara (2006); Ono et al. (2008); Cai et al. (2018, 2020); Doi et al. (2021). Furthermore, SI models have also been developed using SICs Ryu et al. (2004); Shimizu et al. (2013); Ko et al. (2023).
@@ 2.2 Word Order in Simultaneous Interpreting
When dealing with language pairs whose sentence structures are different, including English and Japanese (SVO and head-initial vs. SOV and head-final), reducing the word order difference between the source and the target language is important to minimize the delay.
Murata et al. (2010) segmented source sentences into semantically meaningful units with a maximum length of 4.3 seconds, and translated those units from the viewpoint of SI. He et al. (2015) designed syntactic transformation rules for Japanese-English simultaneous machine translation. By applying the rules to target language sentences (i.e., English), they generated more monotonic translations, while preserving the meaning of source sentences and maintaining the grammaticality of the target language. In English-Japanese SI, Futamata et al. (2020) reordered Japanese sentences to make the word order closer to the original English sentences. They further applied style transfer to make the sentence fluent and obtained sentences close to SI sentences by human interpreters. Han et al. (2021) proposed an algorithm to reorder and refine the target sentences so that the target sentences were aligned largely monotonically. They trained SI models for four language pairs, including English-Japanese. Nakabayashi and Kato (2021) segmented sentences into chunks and created bilingual pairs of the chunks with explicit annotations of context information. The SI model trained on the data translated the source sentences while referencing preceding chunks but connecting chunks naturally remained a challenge. Higashiyama et al. (2023) have constructed a large-scale English Japanese SIC with information of chunk boundaries in source and target sentences and phrases that can be omitted in target sentences. NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset Fukuda et al. (2024), which is similar to Higashiyama et al.’s (2023), is relatively small and intended for the evaluation purpose.
The word order differences among different translation modes have also been investigated. Okamura and Yamada (2023) quantitatively compared the degree to which the word order of the source sentences was maintained and found that SI sentences retained the order better than consecutive interpreting and offline translation sentences. Cai et al. (2020) found syntactic and non-syntactic factors that affect interpreters’ word order decisions through the statistical analyses of an SIC. In this paper, we analyze what makes monotonic translation difficult. While Cai et al. (2020) analyzed actual SI data generated by human simultaneous interpreters, we use the CMT dataset, which is created without time constraints, and in which all information in the source sentences is translated into the target sentences.
@@ 3 Chunk-wise Monotonic Translation
In SI between language pairs with different sentence structures, interpreters are known to segment source sentences into chunks and translate them in the order of the chunks444Reordering may occur within a chunk. Okamura and Yamada (2023). This section describes the details of NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset, which is used in our analyses.
@@ 3.1 Data
NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset consists of 511 pairs of source sentences and their corresponding chunk-wise monotonic translation (CMT) sentences with information of chunk boundaries in source and target sentences. Source (i.e., English) sentences, which were used as the test set in Ko et al. (2023), were segmented following the five rules that reflected interpreter’s strategies555See the original paper or the code for chunking: https://github.com/ahclab/si_chunker based on the syntactic analysis results from spaCy. The source sentences come from 8 TED talks.
| Data | Sum | Per Sent. |
@@
| # chunks | 1,677 | 3.28 |
| # source words | 8,104 | 15.86 |
| # target words | 13,981 | 27.36 |
Translators were provided with source sentences with chunk boundaries and were asked to (1) translate them in the order of the chunks while (2) connecting the chunks naturally and (3) not including the content of the subsequent chunks. It was allowed to (1) repeat, (2) defer, and (3) omit phrases in the source sentences to keep the translation fluent, but was instructed to avoid using the operations with larger number as much as possible. Data statistics are shown in Table 2.
@@ 4 Data Analysis
Fukuda et al. (2024) have examined the quality of CMT sentences through human evaluation but have not analyzed the characteristics of them. We suppose that there are factors that prevent monotonic translation if a phrase in the source sentences is repeated, deferred, and omitted since the CMT sentences were generated without time constraints. We qualitatively and quantitatively analyze the CMT sentences with these operations and reveal the factors.
To better understand the characteristics of CMT sentences, we also compare them with SI sentences from NAIST-SIC and NAIST-SIC-Aligned, and offline translation sentences from the subtitles of TED talks. Since the SI sentences in NAIST-SIC were not aligned at the sentence level, we performed alignment manually. Some source sentences did not match across the datasets, and we excluded those 10 sentences from the analyses. In addition, 25 sentences were not translated in NAIST-SIC666For example, due to time constraints, interpreters may not be able to translate a whole sentence. See Doi et al. (2021)., which were also excluded from the analyses. As a result, 476 sentences were used for our analyses.
@@ 4.1 Annotations
To analyze the characteristics of CMT sentences, we annotated tags to source and CMT sentences. The list of tags is shown in Table 3. Prior to the annotation, we tokenized the English sentences using spaCy777https://spacy.io/ and the Japanese sentences using MeCab Kudo et al. (2004) with unidic. Then, the source and CMT sentences were concatenated with special token [SEP] and annotated using an open-source data labeling tool, doccano888https://github.com/doccano/doccano. Annotation examples are shown in Figure 1.
The first and second authors collaboratively annotated the first 50 examples while discussing. As sufficient agreement was expected, the remainder of the data were annotated by the first author.
4.2 Results of Analyses
@@ 4.2.1 Comparison among Different Translation Modes
We compared the length of sentences in the four datasets. Fukuda et al. (2024) have also conducted similar comparisons based on the number of characters. However, there were variations in spelling and differences in transcription systems (e.g., numbers), and we made comparisons based on the number of words segmented by MeCab.
Table 4 show that CMT sentences were the longest, followed by offline, NAIST-SIC, and NAIST-SIC-Aligned, which matched the results reported in Fukuda et al. (2024). Long translation sentences can pose some problems. As discussed in Fukuda et al. (2024), it may increase the cognitive load on the listeners/readers of the translations. In addition, longer output may cause a delay even though CMT aims to reduce it.
| Dataset | Sum | Per Sent. |
| CMT | 13,508 | 28.38 |
| NAIST-SIC | 8,914 | 18.73 |
| NAIST-SIC-Aligned | 8,072 | 16.96 |
| Offline | 9,907 | 20.81 |
4.2.2 Factors that Make CMT Sentences Longer
To reveal what made the CMT sentences longer, we first analyzed them qualitatively. Our analyses suggest that (1) CMT sentences contain many formulaic expressions for the end of sentences as sentences are segmented into small chunks, and (2) words that are often omitted in Japanese (e.g., pronouns) are explicitly translated since translators were instructed to avoid omitting phrases in the source language, as in the following examples.

### v2

@@ Abstract
This paper analyzes the features of monotonic translations, which follow the word order of the source language, in simultaneous interpreting (SI). Word order differences are one of the biggest challenges in SI, especially for language pairs with significant structural differences like English and Japanese. We analyzed the characteristics of chunk-wise monotonic translation (CMT) sentences using the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset and identified some grammatical structures that make monotonic translation difficult in English-Japanese SI. We further investigated the features of CMT sentences by evaluating the output from the existing speech translation (ST) and simultaneous speech translation (simulST) models on the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset as well as on existing test sets. The results indicate the possibility that the existing SI-based test set underestimates the model performance. The results also suggest that using CMT sentences as references gives higher scores to simulST models than ST models, and that using an offline-based test set to evaluate the simulST models underestimates the model performance.
@@ 1 Introduction
Simultaneous interpreting (SI) is the task of translating speech from a source language into a target language in real time. SI is cognitively demanding, and human simultaneous interpreters employ such strategies as segmentation, summarization, and generalization He et al. (2016). Maintaining word order in a source language is another important strategy, especially for language pairs whose word order differs (e.g., English and Japanese), to shorten delays and reduce cognitive load. Because of these features, SI sentences are different from offline translation sentences, although most automatic SI studies Oda et al. (2014); Ma et al. (2019); Liu et al. (2020); Papi et al. (2023) have used offline translation corpora (e.g., MuST-C; Di Gangi et al., 2019) for both training and evaluatng models due to the limited amount of simultaneous interpretation corpora (SICs).
For English-Japanese language pairs, several SICs have been constructed Toyama et al. (2004); Shimizu et al. (2014); Matsushita et al. (2020); Doi et al. (2021). Based on the NAIST Simultaneous Interpretation Corpus (NAIST-SIC; Doi et al., 2021), Zhao et al. (2024)111The dataset was released in 2023 (see version 3 of the paper). created an automatically-aligned parallel SI dataset: NAIST-SIC-Aligned. Since its sentences are aligned at the sentence level, they can be used for model training. Actually, Ko et al. (2023) and Zhao et al. (2024) trained SI models using SI data from NAIST-SIC-Aligned. Their model performances were evaluated through automatic evaluation metrics such as BLEU Papineni et al. (2002) using a small test set curated based on SI sentences generated by professional human simultaneous interpreters.
Although the scores reported in Ko et al. (2023) and Zhao et al. (2024) were relatively low, the test set used in both studies might have underestimated the model performance. Since human simultaneous interpreters use such strategies as summarization and generalization, phrases that do not affect the main idea are not necessarily translated into the target language. If an SI model generates translations for phrases that a human interpreter did not, the output sentence might not be evaluated properly, even when it is a correct translation.
@@
Fukuda et al. (2024) pointed out the difficulty for SI models to learn which phrases in source speech are less important and advocated constructing SI models that only employ a strategy that maintains the word order in a source language. As a first step, they created the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset222https://dsc-nlp.naist.jp/data/NAIST-SIC/Aligned-Chunk_Mono-EJ/. The source sentences in the test set used in Ko et al. (2023) were automatically segmented into chunks, each of which was translated in a way that did not include the content of subsequent chunks. Unlike in SI sentences by human interpreters, where not all the information in the source sentences is translated, chunk-wise monotonic translation (CMT) sentences333CMT refers to the task of segmenting a source sentence into chunks and translating it in the order of the chunks. A CMT sentence is a target sentence generated through CMT. were translated so that all the information is translated (Table 1)444Precisely, omissions that maintained the fluency of the sentence were allowed. See Section 3.1 for the details about the dataset.. Fukuda et al. (2024) have investigated the quality of the CMT sentences in their dataset through human evaluation, although they have not analyzed its characteristics. Nor have they conducted any evaluation experiments in which model outputs are evaluated on their dataset.
@@
In this paper, we qualitatively and quantitatively analyze CMT sentences in the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset. In the process of generating CMT sentences for the dataset, it was allowed to repeat, defer, and omit phrases in the source sentences to maintain the translation’s fluency. We assume the presence of factors (e.g., syntactic structures) that prevent monotonic translation if phrases were repeated, deferred, or omitted in the CMT sentences since they were translated without time constraints. In addition, we evaluate the output from an existing speech translation (ST) model and two simultaneous speech translation (simulST) models (See 5.2). Both the ST555A ST model generates translations after the utterances are completed. and simulST models are used to investigate the differences in scores when evaluating translations with different characteristics. The contributions of this paper are as follows:
@@
We analyze CMT sentences and show that they tend to be longer than offline translations primarily because of repetition.
@@
We investigate what causes the phrases in source sentences to be repeated, deferred, and omitted and show that most cases occur because of particular grammatical structures. When a phrase in a chunk is a dependent of a phrase in the preceding chunk, the head phrase is typically repeated or deferred.
@@
We evaluate the output from three different models on the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset: (1) an ST model trained on offline data, (2) a simulST model trained on offline data, and (3) a simulST model trained on both offline and SI data. The results suggest that the existing SI-based test set Ko et al. (2023); Zhao et al. (2024) underestimates the model performance. The results also suggest that using CMT sentences as references gives higher scores to simulST models than ST models, while using an offline-based test set for evaluating simulST models underestimates the model performance.
@@ 2.1 Simultaneous Interpretation Corpora
SICs are valuable resources both for developing automatic SI models and analyzing SI’s characteristics. For English-Japanese language pairs, several SICs are publicly available Toyama et al. (2004); Shimizu et al. (2014); Matsushita et al. (2020); Doi et al. (2021), although the amount of such corpora is very limited compared to offline translation corpora.
Using these corpora, SI sentences have been analyzed from various perspectives, such as strategies and interpreting patterns used by interpreters, latency, translation quality, and word order Tohyama and Matsubara (2006); Ono et al. (2008); Cai et al. (2018, 2020); Doi et al. (2021). SI models have also been developed using SICs Ryu et al. (2004); Shimizu et al. (2013); Ko et al. (2023).
@@ 2.2 Word Order in Simultaneous Interpreting
When dealing with language pairs whose sentence structures are different, including English and Japanese (SVO/head-initial vs. SOV/head-final), reducing the word order differences between the source and the target languages is crucial for minimizing delays.
Murata et al. (2010) segmented source sentences into semantically meaningful units with a maximum length of 4.3 seconds and translated those units from an SI viewpoint. He et al. (2015) designed syntactic transformation rules for Japanese-English simultaneous machine translation. By applying the rules to target language sentences (i.e., English), they generated more monotonic translations, while preserving the meaning of source sentences and maintaining the grammaticality of the target language. In English-Japanese SI, Futamata et al. (2020) reordered Japanese sentences to make the word order closer to the original English sentences. They further applied style transfer to increase the fluency and obtained sentences close to SI sentences by human interpreters. Han et al. (2021) proposed an algorithm to reorder and refine the target sentences so that the target sentences were aligned largely monotonically. They trained SI models for four language pairs, including English-Japanese. Nakabayashi and Kato (2021) segmented sentences into chunks and created bilingual pairs of such chunks with explicit annotations of context information. The SI model trained on the data translated the source sentences while referencing the preceding chunks although naturally connecting chunks remained a challenge. Higashiyama et al. (2023) constructed a large-scale English Japanese SIC with the information of chunk boundaries in source and target sentences and phrases that can be omitted in target sentences. The NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset Fukuda et al. (2024), which is similar to Higashiyama et al.’s (2023), is relatively small and intended for the evaluation purposes.
The word order differences among different translation modes have also been investigated. Okamura and Yamada (2023) quantitatively compared the degree to which the word order of the source sentences was maintained and found that SI sentences retained the order better than consecutive interpreting and offline translation sentences. Cai et al. (2020) found syntactic and non-syntactic factors that affect interpreters’ word order decisions through the statistical analyses of an SIC. In this paper, we analyze what makes monotonic translation difficult. While Cai et al. (2020) analyzed the actual SI data generated by human simultaneous interpreters, we use CMT sentences, which were generated without time constraints, and in which all the information in the source sentences is translated into target sentences.
@@ 3 Chunk-wise Monotonic Translation
In SI between language pairs with different sentence structures, interpreters segment source sentences into chunks and translate them from chunk to chunk666Reordering may occur within a chunk. Okamura and Yamada (2023). This section describes the details of the NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset, which is used in our analyses.
@@ 3.1 Data
The NAIST English-to-Japanese Chunk-wise Monotonic Translation Evaluation Dataset consists of 511 pairs of source sentences and their corresponding chunk-wise monotonic translation (CMT) sentences with information of chunk boundaries in the source and target sentences. Source (i.e., English) sentences, which were used as the test set in Ko et al. (2023), were segmented following the five rules that reflected the interpreter’s strategies777See the original paper or the code for chunking: https://github.com/ahclab/si_chunker based on the syntactic analysis results from spaCy. The source sentences come from eight TED talks.
| Data | Sum | Per Sent.SD |
@@
| # chunks | 1,677 | 3.282.12 |
| # source words | 8,104 | 15.8610.16 |
| # target words | 13,981 | 27.3618.55 |
Translators were provided with source sentences with chunk boundaries and asked to (1) translate them in the order of the chunks while (2) naturally connecting the chunks and (3) not including the content of subsequent chunks. They were allowed to (1) repeat, (2) defer, and (3) omit phrases in the source sentences to keep translation fluency, although they were instructed to minimize their use of the operations with larger number as much as possible (e.g., defer should not be used when repeat can handle the situation) . Data statistics are shown in Table 2.
@@ 4 Data Analysis
Fukuda et al. (2024) have examined the quality of CMT sentences through human evaluation but have not analyzed the characteristics of them. We suppose that factors exist that prevent monotonic translation if a phrase in the source sentences is repeated, deferred, or omitted since the CMT sentences were generated without time constraints. Therefore we qualitatively and quantitatively analyze the CMT sentences with these operations and reveal such factors.
To better understand the characteristics of the CMT sentences, we also compare them with SI sentences from NAIST-SIC and NAIST-SIC-Aligned as well as offline translation sentences from the subtitles of TED talks. Since the SI sentences in NAIST-SIC were not aligned at the sentence level, we manually align them. Some source sentences did not match across the datasets, and we excluded those ten sentences from the analyses. In addition, 25 sentences were not translated in NAIST-SIC888For example, due to time constraints, interpreters might have been unable to translate a whole sentence. See Doi et al. (2021)., which were also excluded from the analyses. As a result, 476 sentences were used for our analyses.
@@ 4.1 Annotations
To analyze the characteristics of the CMT sentences, we annotated tags to the source and CMT sentences. The list of tags is shown in Table 3. Prior to the annotations, we tokenized the English sentences using spaCy999https://spacy.io/ and the Japanese sentences using MeCab Kudo et al. (2004) with unidic. Then, we concatenated the source and CMT sentences with a special token [SEP] and annotated them using an open-source data labeling tool, doccano.101010https://github.com/doccano/doccano
We identified spans (i.e., words or phrases) that are repeated, deferred, or omitted and annotated the span tags. In addition, we annotated ahead, add, and error tags for analyses of problematic translations. The corresponding span tags in the source and CMT sentences were associated using relation tags. Annotation examples are shown in Figure 1.
The first and second authors collaboratively annotated the first 50 examples while discussing their decisions. Since sufficient agreement was assumed, the remainder of the data were just annotated by the first author.
4.2 Analysis Results
@@ 4.2.1 Comparison among Different Translation Modes
We compared the sentence lengths of the four datasets. Fukuda et al. (2024) also conducted similar comparisons based on the number of characters. However, since variations in spelling and differences in transcription systems (e.g., numbers) were found, we made comparisons based on the number of words segmented by MeCab.
Table 4 shows that the CMT sentences were the longest, followed by offline, NAIST-SIC, and NAIST-SIC-Aligned. These results matched those reported in Fukuda et al. (2024). Long translated sentences can pose some problems. As discussed in Fukuda et al. (2024), the lenght may increase the cognitive load on the listeners/readers of the translations. In addition, longer output may cause a delay even though CMT aims to reduce it.
| Dataset | Sum | Per Sent.SD |
| CMT | 13,508 | 28.3818.66 |
| NAIST-SIC | 8,914 | 18.7312.08 |
| NAIST-SIC-Aligned | 8,072 | 16.9611.52 |
| Offline | 9,907 | 20.8112.62 |
4.2.2 Factors that Lengthen CMT Sentences
To reveal what factor lengthened the CMT sentences, we first analyzed them qualitatively. Our analyses suggest that (1) CMT sentences contain many formulaic expressions for the end of sentences as they are segmented into small chunks, and (2) the words that are often omitted in Japanese (e.g., pronouns) are explicitly translated since translators were instructed to avoid omitting phrases in the source language, as in the following examples: